OpenStudy (linda):

if the 3rd term of an arithmetic sequence is 34 and the 6th term is 55, find the common difference, the 1st term of the sequence, a formula for the nth term of the sequence, and the sum of the first 50 terms. can anyone help? much appreciated!!

OpenStudy (anonymous):

a=20, d=7, a(n)=7n+13, S(50)=9575

OpenStudy (amistre64):

A3 = 34 A6 = 55 34 + 3n = 55 (55-34)/3 = n n = 7

OpenStudy (amistre64):

A{x} A{3} = 7(x+2) -1 A{6} = 7(x+2) -1 A{x} = 7(x+2) -1 A{50} = 7(50+2) -1 A{50} = 7(52) -1 A(50) = 364 -1 = 363 is what I get

OpenStudy (amistre64):

if we use xaviers; it matches: A{50} = 7(50) + 13 = 363

OpenStudy (amistre64):

and the 'sum; of the first 50 terms eh .......

OpenStudy (linda):

what do you mean by A?

OpenStudy (amistre64):

A is just a name for the function; i could have called it Bob ... but A just came to mind

OpenStudy (amistre64):

the sum is prolly a multiple of 7 and a adjustment for it not being actually '7'